A vending machine will accept nickels and dimes only. The attendant checks the machine and finds a total of 42 coins whose value is $3.40. How many coins of each type were there?
Answer by checkley77(12569)

D+N=42 OR D=42-N
.10D+.05N=3.40
.10(42-N)+.05N=3.40
4.20-.10N+.05N=3.40
-.05N=3.40-4.20
-.05N=-.80
N=-.80/-.05
N=16 NUMBER OF NICKELS.
42-16=26 NUMBER OF DIMES.
PROOF:
.10*26+.05*16=3.40
2.6+.80=3.40
3.40=3.40